I am really confused by this question
find a polynormal for the perimeter and for the area x+2 length and x the width
the perimeter is___________do not factor
the area is______________do not factor

Answer by oberobic(2304)

L = length
W = width
L = x+2
W = x
.
Perimeter of rectangle = 2(L+W)
2(L+W) = 2(x+2+x)
= 2(2x+2)
= 4x +4
.
Area of rectangle = L*W
L*W = (x+2)*x
= x^2 +2x
.
Done